# An explicit form of Ingham’s zero density estimate

### v1

@@ Abstract.
Ingham (1940) proved that $N(\sigma,T)\ll T^{3(1-\sigma)/(2-\sigma)}\log^{5}{T}$, where $N(\sigma,T)$ counts the number of the non-trivial zeros $\rho$ of the Riemann zeta-function with $\Re\{\rho\}\geq\sigma\geq 1/2$ and $0<\Im\{\rho\}\leq T$. We provide an explicit version of this result with the exponent $3$ of the logarithmic factor. In addition, we also provide an explicit estimate with asymptotically correct main term for the fourth power moment of the Riemann zeta-function on the critical line.
@@ 1. Introduction
Let $\zeta(s)$ be the Riemann zeta-function and let $\rho=\beta+\mathrm{i}\gamma$ be its non-trivial zero, that is, $\zeta(\rho)=0$ with $0\leq\beta\leq 1$ and $\gamma\in\mathbb{R}$. Let $N(T)$ be the number (counted with multiplicities) of all non-trivial zeros with $0<\gamma\leq T$, while for $\sigma\geq 0$ and $T>0$ define $N(\sigma,T)\mathrel{\mathop{:}}=\#\left\{\rho\colon\zeta(\rho)=0,\beta\geq\sigma,0<\gamma\leq T\right\}$, where the zeros are again counted with multiplicities. Trivially, the function $N(\sigma,T)$ is decreasing in $\sigma$ and increasing in $T$, $N(\sigma,T)\leq N(T)\sim\frac{1}{2\pi}T\log{T}$ for $20\leq\sigma\leq 1/$ 2, $N(\sigma,T)\leq\frac{1}{2}N(T)$ for $11/2<\sigma<$ 1 and $N(\sigma,T)=0$ for $\sigma\geq 1$, while the Riemann Hypothesis (RH) asserts that $N(\sigma,T)=N(T)$ for $20\leq\sigma\leq 1/$ 2 and $N(\sigma,T)=0$ for $\sigma>1/2$. Since RH has been verified up to $H_{\mathrm{RH}}\mathrel{\mathop{:}}=3\cdot 10^{12}$ by Platt and Trudgian [PT21], we have $N(\sigma,T)=0$ for $\sigma>1/2$ and $T\leq H_{\mathrm{RH}}$. Consequently, we can restrict ourselves to the range of $1/2\leq\sigma\leq 1$ and $T\geq H_{\mathrm{RH}}$ throughout this paper.
@@
with $A(\sigma)\leq 1/(1-\sigma)$, $B\geq 0$, $\varepsilon>0$ and the implied constants being uniform in $\sigma$ and $T$, are of great interest in the distribution theory of prime numbers [Ivi03, Sta26]. The Lindelöf Hypothesis (LH), which is believed to be a weaker statement than RH, implies $A(\sigma)\leq 2$ in the expression on the right-hand side of (1) for all $\sigma\in[1/2,1]$, and this is known as the density hypothesis [Ivi03, p. 47]. Bourgain [Bou00] showed that the density hypothesis holds for $\sigma\geq 25/32=0.78125$, and substantial progress has been made in obtaining much better upper bounds for $A(\sigma)$ in this region, see [TTY25, Table 2 on p. 33] for the current best results. On the other hand, region $0.761/2\leq\sigma<19/25=$ 0.76 is dominated by two results, namely that
@@
In order to obtain effective and unconditional results that are related to the distribution of prime numbers, explicit versions of (1) are often required, see [FKS23, JY23, CHJ25]. Recently, several relevant results have appeared in the literature, see [Kad13, KLN18, Bel25, Cho24] for example. The aim of this paper is to provide the first explicit version of (2), see Corollary 1, where we also made an improvement over the logarithmic factor.
@@
$N(\sigma,T)\leq\mathfrak{B}_{1}T^{\frac{3(1-\sigma)}{2-\sigma}}\log^{3}{T}+\mathfrak{B}_{2}\log^{2}{T}+\mathfrak{B}_{3}\log{T}$ (3)
@@
The intervals $[\sigma_{1},\sigma_{2}]$ in Table 1 are elements of the set $\left\{[n/32,(n+1)/32]\colon n\in\{16,\ldots,31\}\right\}$. This choice is justified by the fact that (3) is asymptotically better than $N(\sigma,T)\ll T^{8(1-\sigma)/3}\log^{5}{T}$ for $\sigma\in[1/2,0.875]$ and the latter zero density estimate [^4] We are comparing our result to [KLN18], where a variant of such a zero density estimate was proved. is asymptotically better than the trivial one for $\sigma\in(0.625,1)$. Corollary 1 implies that
$N(\sigma,T)\leq 8.604\cdot T^{\frac{3(1-\sigma)}{2-\sigma}}\log^{3}{T}+9.461\log^{2}{T}+167.8\log{T},\quad\sigma\in[0.500,0.625],\;T\geq 3\cdot 10^{12},$
$N(\sigma,T)\leq 22.44\cdot T^{\frac{3(1-\sigma)}{2-\sigma}}\log^{3}{T}+8.290\log^{2}{T}+147.0\log{T},\quad\sigma\in[0.625,0.875],\;T\geq 3\cdot 10^{12}.$
The above estimates improve [Cho24, Theorem 1.1 and Table 1] for $\sigma\in[0.5,0.667]$ and the latter one improves [KLN18, Theorem 1.1 and Table 1] for $\sigma\in[0.625,0.875]$, for $T\geq 3\cdot 10^{12}$.
$\sigma_{1}$ $\sigma_{2}$ $d$ $\kappa$ $\delta$ $\log_{10}A_{0}$ $\mathfrak{B}_{1}$ $\mathfrak{B}_{2}$ $\mathfrak{B}_{3}$
$0.50000$ $0.53125$ $0.28$ $0.36$ $0.012$ $10.3$ $5.360$ $9.461$ $167.8$
$0.53125$ $0.56250$ $0.28$ $0.34$ $0.025$ $10.4$ $6.380$ $9.168$ $162.6$
$0.56250$ $0.59375$ $0.28$ $0.32$ $0.038$ $10.4$ $7.453$ $8.875$ $157.4$
$0.59375$ $0.62500$ $0.28$ $0.30$ $0.051$ $10.4$ $8.604$ $8.582$ $152.2$
$0.62500$ $0.65625$ $0.28$ $0.28$ $0.066$ $10.4$ $9.848$ $8.290$ $147.0$
$0.65625$ $0.68750$ $0.28$ $0.26$ $0.080$ $10.4$ $11.20$ $7.997$ $141.9$
$0.68750$ $0.71875$ $0.28$ $0.24$ $0.096$ $10.6$ $12.65$ $7.704$ $136.7$
$0.71875$ $0.75000$ $0.28$ $0.22$ $0.112$ $10.6$ $14.23$ $7.411$ $131.5$
$0.75000$ $0.78125$ $0.28$ $0.20$ $0.129$ $10.6$ $15.95$ $7.118$ $126.3$
$0.78125$ $0.81250$ $0.28$ $0.18$ $0.147$ $10.6$ $17.84$ $6.826$ $121.1$
$0.81250$ $0.84375$ $0.28$ $0.16$ $0.166$ $10.6$ $19.97$ $6.533$ $115.9$
$0.84375$ $0.87500$ $0.28$ $0.15$ $0.187$ $10.6$ $22.44$ $6.240$ $110.7$
$0.87500$ $0.90625$ $0.28$ $0.13$ $0.209$ $10.8$ $25.51$ $5.947$ $105.5$
$0.90625$ $0.93750$ $0.29$ $0.11$ $0.234$ $10.8$ $29.71$ $5.463$ $96.87$
$0.93750$ $0.96875$ $0.29$ $0.09$ $0.261$ $10.8$ $36.09$ $5.180$ $91.86$
$0.96875$ $1.00000$ $0.29$ $0.07$ $0.292$ $10.8$ $46.92$ $4.897$ $86.84$
@@
is believed to be true. Unconditionally, the upper bound in (5) is known to hold for $0<k\leq 2$, see [HRS19], while the lower bound is true for all $k\geq 1$, see [RS13]. On RH, the upper bound in (5) also holds for $k>2$, see [Har13], while the lower bound is also true for $10<k<$ 1, see [HB81]. The random matrix theory predicts the asymptotic form of (5) for $k\in\mathbb{N}$, that is,
@@
for $\varepsilon>0$, where $p_{k}(x)$ is a polynomial of degree $k^{2}$ whose coefficients can be effectively computed, see [CFK+05]. Polynomials $p_{k}(x)$ from (6) are known only for $k\in\{1,2\}$: $p_{1}(x)=x+2\gamma-1-\log{(2\pi)}$ where $\gamma$ is the Euler–Mascheroni constant is due to Ingham [Ing26], and $p_{2}(x)=\frac{1}{2\pi^{2}}x^{4}+\sum_{k=1}^{4}a_{4-k}x^{4-k}$ is due to Heath-Brown [HB79], where his [^6] His error term is of order $T^{7/8+\varepsilon}$. This was improved in [IM95], see also [PT25], to $T^{2/3+\varepsilon}$ by using methods from the spectral theory of the Riemann zeta-function. values for $a_{k}$, $k\in\{0,1,2,3\}$, coincide with those predicted by the random matrix theory. It should be mentioned that Ingham proved in the same paper [Ing26] that $p_{2}(x)=\frac{1}{2\pi^{2}}x^{4}+O(x^{3})$, i.e.,
@@
thus providing an asymptotic formula for $\mathcal{M}_{2}(T)$. Recently, explicit estimates for the second power moment $\mathcal{M}_{1}(T)$ appeared in [DZA22] and [SS23] that are of the form (6) but with better error terms. To the best of the authors knowledge, no explicit estimates for the fourth power moment $\mathcal{M}_{2}(T)$ have appeared in the literature. In this paper, we are providing them that are of the form $\ll T\log^{4}{T}$ and (7), see Corollary 2.
@@
$-\left(48.942+\frac{10^{5}\log{T}}{\pi^{2}T}\right)T\log^{3}{\frac{T}{2}}\leq\mathcal{M}_{2}(T)-\frac{1}{2\pi^{2}}T\log^{4}{\frac{T}{2}}\leq 48.801\cdot T\log^{3}{\frac{T}{2}}+3.0592\cdot 10^{10}$ (9)
for $T\geq 10^{5}$.
@@
$\mathcal{M}_{\frac{3}{2}}(T)\leq\left(\frac{1}{2\pi^{2}}+\frac{241.03}{\log{T_{0}}}+\frac{132.1}{\log^{2}{T_{0}}}\right)^{\frac{1}{2}}T\log^{\frac{5}{2}}{T}$
for $T\geq T_{0}\geq 10^{5}$.
@@
The outline of this paper is as follows. In Section 2 we state the main results, namely Theorem 1 for the zero density estimate and Theorem 2 for the fourth power moment estimate, and prove the above corollaries. In Section 3 we prove Gabriel’s convexity theorem, together with two applications (Corollaries 6 and 7), and in Section 4 we prove Theorem 1. Section 5 is dedicated to the proof of Theorem 2.
@@ Acknowledgements
The authors thank Andrew Yang and Tim Trudgian for their interest in the work.

### v2

@@ Abstract.
Ingham (1940) proved that $N(\sigma,T)\ll T^{3(1-\sigma)/(2-\sigma)}\log^{5}{T}$, where $N(\sigma,T)$ counts the number of the non-trivial zeros $\rho$ of the Riemann zeta-function with $\Re\{\rho\}\geq\sigma\geq 1/2$ and $0<\Im\{\rho\}\leq T$. We provide an explicit version of this result with the exponent $(7-5\sigma)/(2-\sigma)$ of the logarithmic factor. In addition, we also provide an explicit estimate with asymptotically correct main term for the fourth power moment of the Riemann zeta-function on the critical line.
@@ 1. Introduction
Let $\zeta(s)$ be the Riemann zeta-function and let $\rho=\beta+\mathrm{i}\gamma$ be its non-trivial zero, that is, $\zeta(\rho)=0$ with $0\leq\beta\leq 1$ and $\gamma\in\mathbb{R}$. Let $N(T)$ be the number (counted with multiplicities) of all non-trivial zeros with $0<\gamma\leq T$, while for $\sigma\geq 0$ and $T>0$ define $N(\sigma,T)\mathrel{\mathop{:}}=\#\left\{\rho\colon\zeta(\rho)=0,\beta\geq\sigma,0<\gamma\leq T\right\}$, where the zeros are again counted with multiplicities. Trivially, the function $N(\sigma,T)$ is decreasing in $\sigma$ and increasing in $T$, $N(\sigma,T)\leq N(T)\sim\frac{1}{2\pi}T\log{T}$ for $0\leq\sigma\leq 1/2$, $N(\sigma,T)\leq\frac{1}{2}N(T)$ for $1/2<\sigma<1$ and $N(\sigma,T)=0$ for $\sigma\geq 1$, while the Riemann Hypothesis (RH) asserts that $N(\sigma,T)=N(T)$ for $0\leq\sigma\leq 1/2$ and $N(\sigma,T)=0$ for $\sigma>1/2$. Since RH has been verified up to $H_{\mathrm{RH}}\mathrel{\mathop{:}}=3\cdot 10^{12}$ by Platt and Trudgian [PT21], we have $N(\sigma,T)=0$ for $\sigma>1/2$ and $T\leq H_{\mathrm{RH}}$. Consequently, we can restrict ourselves to the range of $1/2\leq\sigma\leq 1$ and $T\geq H_{\mathrm{RH}}$ throughout this paper.
@@
with $A(\sigma)\leq 1/(1-\sigma)$, $B\geq 0$, $\varepsilon>0$ and the implied constants being uniform in $\sigma$ and $T$, are of great interest in the distribution theory of prime numbers [Ivi03, Sta26]. The Lindelöf Hypothesis (LH), which is believed to be a weaker statement than RH, implies $A(\sigma)\leq 2$ in the expression on the right-hand side of (1) for all [^2] It is known that LH guarantees even $A(\sigma)=0$ for $\sigma\geq\frac{3}{4}+\delta$ and $\delta>0$ fixed. $\sigma\geq 1/2$, and this is known as the density hypothesis [Ivi03, p. 47]. Bourgain [Bou00] showed that the density hypothesis holds for $\sigma\geq 25/32=0.78125$, and substantial progress has been made in obtaining much better upper bounds for $A(\sigma)$ in this region, see [TTY25, Table 2 on p. 33] for the current best results. On the other hand, region $1/2\leq\sigma<19/25=0.76$ is dominated by two results, namely that
@@
In order to obtain effective and unconditional results that are related to the distribution of prime numbers, explicit versions of (1) are often required, see [FKS23, JY23, CHJ23, CHJ25]. Recently, several relevant results have appeared in the literature, see [Kad13, KLN18, Bel25, Cho24] for example. The aim of this paper is to provide the first explicit version of (2), see Corollary 1, where we also made an improvement over the logarithmic factor.
@@
$N(\sigma,T)\leq\mathfrak{B}_{1}T^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{T}\right)^{\frac{7-5\sigma}{2-\sigma}}+\mathfrak{B}_{2}\log^{2}{T}+\mathfrak{B}_{3}\log{T}$ (3)
@@
Note that $(7-5\sigma)/(2-\sigma)$ is a decreasing function in $\sigma\in[1/2,1]$ that has the range $[2,3]$ there. The intervals $\sigma_{1},\sigma_{2}]$ in Table 1 are elements of the set $\left\{[n/32,(n+1)/32]\colon n\in\{16,\ldots,31\}\right\}$. This choice is justified by the fact that (3) is asymptotically better than $N(\sigma,T)\ll T^{8(1-\sigma)/3}\log^{5}{T}$ for $\sigma\in[1/2,0.875]$ and the latter zero density estimate [^5] We are comparing our result to an improved version (see Section 6) of [KLN18, Theorem 1.1 and Table 1], see also [CHJ23, Lemma 7.1], where a variant of such a zero density estimate was proved. is asymptotically better than the trivial one for $\sigma\in(0.625,1)$. Corollary 1 implies that
$N(\sigma,T)\leq 8.185\cdot T^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{T}\right)^{\frac{7-5\sigma}{2-\sigma}}+9.461\log^{2}{T}+167.8\log{T},\quad\sigma\in[0.500,0.625],\;T\geq 3\cdot 10^{12},$
$N(\sigma,T)\leq 21.77\cdot T^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{T}\right)^{\frac{7-5\sigma}{2-\sigma}}+8.290\log^{2}{T}+147.0\log{T},\quad\sigma\in[0.625,0.875],\;T\geq 3\cdot 10^{12}.$
The above estimates improve [Cho24, Theorem 1.1 and Table 1] for $\sigma\in[0.5,0.667]$, and the latter improves (119) for $\sigma\in[0.625,0.8575]$ while it is also better for $\sigma\in(0.8575,0.875]$ and $T\geq 10^{61.3}$.
$\sigma_{1}$ $\sigma_{2}$ $\mathfrak{B}_{1}$ $\mathfrak{B}_{2}$ $\mathfrak{B}_{3}$ $d$ $\kappa$ $\delta$ $\log_{10}A_{0}$
$0.50000$ $0.53125$ $5.069$ $9.461$ $167.8$ $0.28$ $0.34$ $0.012$ $11.25$
$0.53125$ $0.56250$ $6.046$ $9.168$ $162.6$ $0.28$ $0.32$ $0.025$ $11.26$
$0.56250$ $0.59375$ $7.077$ $8.875$ $157.4$ $0.28$ $0.30$ $0.038$ $11.27$
$0.59375$ $0.62500$ $8.185$ $8.582$ $152.2$ $0.28$ $0.29$ $0.052$ $11.27$
$0.62500$ $0.65625$ $9.387$ $8.290$ $147.0$ $0.28$ $0.27$ $0.066$ $11.28$
$0.65625$ $0.68750$ $10.70$ $7.997$ $141.9$ $0.28$ $0.25$ $0.081$ $11.29$
$0.68750$ $0.71875$ $12.11$ $7.704$ $136.7$ $0.28$ $0.23$ $0.097$ $11.30$
$0.71875$ $0.75000$ $13.66$ $7.411$ $131.5$ $0.28$ $0.21$ $0.114$ $11.31$
$0.75000$ $0.78125$ $15.35$ $7.118$ $126.3$ $0.28$ $0.20$ $0.132$ $11.32$
$0.78125$ $0.81250$ $17.21$ $6.826$ $121.1$ $0.28$ $0.18$ $0.150$ $11.33$
$0.81250$ $0.84375$ $19.31$ $6.533$ $115.9$ $0.28$ $0.16$ $0.170$ $11.35$
$0.84375$ $0.87500$ $21.77$ $6.240$ $110.7$ $0.28$ $0.14$ $0.192$ $11.36$
$0.87500$ $0.90625$ $24.83$ $5.745$ $101.9$ $0.29$ $0.13$ $0.215$ $11.37$
$0.90625$ $0.93750$ $29.01$ $5.463$ $96.87$ $0.29$ $0.11$ $0.241$ $11.40$
$0.93750$ $0.96875$ $35.35$ $5.180$ $91.86$ $0.29$ $0.09$ $0.270$ $11.42$
$0.96875$ $1.00000$ $46.06$ $4.897$ $86.84$ $0.29$ $0.07$ $0.303$ $11.46$
@@
is believed to be true. Unconditionally, the upper bound in (5) is known to hold for $0<k\leq 2$, see [HRS19], while the lower bound is true for all $k\geq 1$, see [RS13]. On RH, the upper bound in (5) also holds for $k>2$, see [Har13], while the lower bound is also true for $0<k<1$, see [HB81]. Random matrix theory predicts the asymptotic form of (5) for $k\in\mathbb{N}$, that is,
@@
for $\varepsilon>0$, where $p_{k}(x)$ is a polynomial of degree $k^{2}$ whose coefficients can be effectively computed, see [CFK+05]. Polynomials $p_{k}(x)$ from (6) are known only for $k\in\{1,2\}$: $p_{1}(x)=x+2\gamma-1-\log{(2\pi)}$ where $\gamma$ is the Euler–Mascheroni constant, is due to Ingham [Ing26], and $p_{2}(x)=\frac{1}{2\pi^{2}}x^{4}+\sum_{k=1}^{4}a_{4-k}x^{4-k}$ is due to Heath-Brown [HB79], where his [^7] His error term is of order $T^{7/8+\varepsilon}$. This was improved in [IM95], see also [PT25], to $T^{2/3+\varepsilon}$ by using methods from the spectral theory of the Riemann zeta-function. values for $a_{k}$, $k\in\{0,1,2,3\}$, coincide with those predicted by the random matrix theory. It should be mentioned that Ingham proved in the same paper [Ing26] that $p_{2}(x)=\frac{1}{2\pi^{2}}x^{4}+O(x^{3})$, i.e.,
@@
thus providing an asymptotic formula for $\mathcal{M}_{2}(T)$. Recently, explicit estimates for the second power moment $\mathcal{M}_{1}(T)$ appeared in [DZA22] and [SS23] that are of the form (6) but with better error terms. To the best of the authors’ knowledge, no explicit estimates for the fourth power moment $\mathcal{M}_{2}(T)$ have appeared in the literature. In this paper, we are providing them in Corollary 2.
@@
$\left(\mathfrak{M}_{1}(T_{0})+\frac{2\log{2}}{\pi^{2}}+\frac{T_{0}\log{\frac{T}{2}}}{\pi^{2}T}\right)T\log^{3}{\frac{T}{2}}\leq\mathcal{M}_{2}(T)-\frac{1}{2\pi^{2}}T\log^{4}{\frac{T}{2}}\leq\mathfrak{M}_{1}(T_{0})T\log^{3}{\frac{T}{2}}+\mathfrak{M}_{2}(T_{0})$ (9)
for $T\geq T_{0}$, where the values for $\mathfrak{M}_{1}(T_{0})$ and $\mathfrak{M}_{2}(T_{0})$ are given by Table 2.
@@
$T_{0}$ $\mathfrak{M}_{1}(T_{0})$ $\mathfrak{M}_{2}(T_{0})$ $\mathcal{M}_{2}(T_{0})\leq$ $\mathcal{M}_{2}(T_{0},2T_{0})\leq$
$10^{4}$ $59.300$ $1.3796\cdot 10^{9}$ $3.8095\cdot 10^{8}$ $9.9857\cdot 10^{8}$
$10^{5}$ $48.801$ $2.8051\cdot 10^{10}$ $8.7748\cdot 10^{9}$ $1.9276\cdot 10^{10}$
$10^{6}$ $43.230$ $4.1722\cdot 10^{11}$ $1.3688\cdot 10^{11}$ $2.8034\cdot 10^{11}$
$10^{7}$ $39.720$ $5.8170\cdot 10^{12}$ $1.8478\cdot 10^{12}$ $3.9692\cdot 10^{12}$
$10^{8}$ $37.550$ $7.9793\cdot 10^{13}$ $2.5002\cdot 10^{13}$ $5.4791\cdot 10^{13}$
$10^{9}$ $36.211$ $1.0629\cdot 10^{15}$ $3.3317\cdot 10^{14}$ $7.2965\cdot 10^{14}$
$10^{10}$ $34.842$ $1.3897\cdot 10^{16}$ $4.3883\cdot 10^{15}$ $9.5079\cdot 10^{15}$
$10^{15}$ $31.170$ $4.4328\cdot 10^{21}$ $1.4174\cdot 10^{21}$ $3.0154\cdot 10^{21}$
$10^{20}$ $29.364$ $9.6675\cdot 10^{26}$ $3.1233\cdot 10^{26}$ $6.5442\cdot 10^{26}$
Table 2. The values for $\mathfrak{M}_{1}(T_{0})$ and $\mathfrak{M}_{2}(T_{0})$ from Corollaries 2 and 3, as well as upper bounds for $\mathcal{M}_{2}(T_{0})$ and $\mathcal{M}_{2}(T_{0},2T_{0})$ that are used in the proof of Corollary 2.
@@
$\mathcal{M}_{\frac{3}{2}}(T)\leq\left(\frac{1}{2\pi^{2}}+\frac{\mathfrak{M}_{1}(T_{0})}{\log{T}}+\frac{\mathfrak{M}_{2}(T_{0})}{T\log^{4}{T}}\right)^{\frac{1}{2}}T\log^{\frac{5}{2}}{T}$
for $T\geq T_{0}\geq 10^{6}$, where the values for $\mathfrak{M}_{1}(T_{0})$ and $\mathfrak{M}_{2}(T_{0})$ are given by Table 2.
@@
The outline of this paper is as follows. In Section 2 we state the main results, namely Theorem 1 for the zero density estimate and Theorem 2 for the fourth power moment estimate, and prove the above corollaries. In Section 3 we prove Gabriel’s convexity theorem, together with two applications (Corollaries 6 and 7), and in Section 4 we prove Theorem 1. Section 5 is dedicated to the proof of Theorem 2, while in Section 6 we state an improved explicit version of Ingham’s earlier zero density estimate.
@@ Acknowledgements
The authors thank Andrew Yang, Tim Trudgian, Olivier Ramaré and Daniel Johnston for their interest in the work.